OpenStudy (anonymous):

How do I write an equation and determine the 8th term: 7, 7.5, 7.75, 7.875...?

OpenStudy (nottim):

Dunno if this helps, but X/Y chart?

OpenStudy (turingtest):

we are adding 0.5,0.25.,0.125 so the next term should be 7.875+(0.125/2)=7.9375

OpenStudy (anonymous):

yup... but what kind of sequence is it? geometric or using the recursion formula?

OpenStudy (anonymous):

or arithmetic?

OpenStudy (turingtest):

sorry for the typos...

OpenStudy (anonymous):

because i have to show my work on how I got the answer.

OpenStudy (turingtest):

umm... I forgot the definitions, but it's not geometric. \[n_0=7\]\[n_1=7+(0.5)^1\]\[n_2=n_1+(0.5)^2\]in general\[n_i=n_{i-1}+(1/2)^i=7+1/2^1+1/2^2+...+1/2^{i-1}+1/2^i\]

OpenStudy (anonymous):

ohhh!! It's recursion formula? Is it like when t1 = x; tn-1 = tn+1 - tn... etc?

OpenStudy (turingtest):

You can write it recursively as I did at first, but the last formula I gave is not recursive eighth term:\[n_7=7+1/2^1+1/2^2+1/2^3+1/2^4+1/2^5+1/2^6+1/2^7\]

OpenStudy (anonymous):

ohh..

OpenStudy (turingtest):

the last way I wrote the formula that is not recursive is \[n_i=n_0+1/2^1+1/2^2+...+1/2^{i-1}+1/2^i\]the recursive way is\[n_i=n_{i-1}+1/2^i\](note the two are equivalent)

OpenStudy (anonymous):

yup. Oouu, okay! Thank you!!